OpenStudy (anonymous):

Let y represent the total cost of publishing a book (in dollars). Let x represent the number of copies of the book printed. Suppose that x and x are related by the equation . what is the change in the total cost for each book printed? what is the cost to get started (before any books are printed)?

OpenStudy (anonymous):

equation y=1250+25x

OpenStudy (ddcamp):

Let's say you don't print any books, and just get it set up (part 2): x (books printed) = 0, so y (cost) = 1250 + 25(0) = 1250

OpenStudy (anonymous):

so 1250 dollars for the second one?

OpenStudy (ddcamp):

Yes.

OpenStudy (anonymous):

what is the first one then

OpenStudy (ddcamp):

Let's say you print just 1 book. The total cost would be: y = 1250 + 25(1) = 1275 What's the difference in cost between printing 1 book and 0 books?

OpenStudy (anonymous):

25 dollars

OpenStudy (ddcamp):

Yup. That is the change per book printed, the answer to part 1.

OpenStudy (anonymous):

thanks